OpenStudy (anonymous):

Find all pairs (x,y) that satisfy both of the following equations simultaneously. 2x+3y=5 and x-2y=8

OpenStudy (anonymous):

2x +3y = 5 x - 2y = 8 mult by -2 -2x + 4y = -16 add 1st and 3rd eqns 7y = -11 y = - 11/7 2x -33/7 = 5 x = [5 + 33/7]/2 = other part of answer